OpenStudy (anonymous):

find the x-intercepts (show work) (x+1)^2-16 I know one is 3 and the other is -5 but i dont know how to get -5

OpenStudy (callisto):

Put (x+1)^2-16 =0 (x+1)^2-16 = 0 (x+1)^2 = 16 x+1 = 4 or x+1 = -4 <- solve these 2 equations.

OpenStudy (anonymous):

thank you

OpenStudy (callisto):

You're welcome! And welcome to OpenStudy too!